### v1

# Jailbreak Attacks and Defenses against Multimodal Generative Models: A Survey

Abstract
The rapid evolution of multimodal foundation models has led to significant advancements in cross-modal understanding and generation across diverse modalities, including text, images, audio, and video. However, these models remain susceptible to jailbreak attacks, which can bypass built-in safety mechanisms and induce the production of potentially harmful content. Consequently, understanding the methods of jailbreak attacks and existing defense mechanisms is essential to ensure the safe deployment of multimodal generative models in real-world scenarios, particularly in security-sensitive applications. To provide comprehensive insight into this topic, this survey reviews jailbreak and defense in multimodal generative models. First, given the generalized lifecycle of multimodal jailbreak, we systematically explore attacks and corresponding defense strategies across four levels: input, encoder, generator, and output. Based on this analysis, we present a detailed taxonomy of attack methods, defense mechanisms, and evaluation frameworks specific to multimodal generative models. Additionally, we cover a wide range of input-output configurations, including modalities such as Any-to-Text, Any-to-Vision, and Any-to-Any within generative systems. Finally, we highlight current research challenges and propose potential directions for future research. The open-source repository corresponding to this work can be found at https://github.com/liuxuannan/Awesome-Multimodal-Jailbreak.
Index Terms:
Jailbreak, Multimodal, Generative Model.
I Introduction
In recent years, multimodal generative models have made significant advancements in both understanding and generation [1, 2]. For multimodal understanding, Multimodal Large Language Models (MLLMs) [3, 4, 5] have demonstrated notable capabilities in Any-to-Text comprehension, excelling in tasks such as visual, audio, and video question-answering [6, 7, 8]. For multimodal generation, denoising diffusion probabilistic models (DDPMs) [9] have achieved impressive performance in Any-to-Vision generation [10, 11, 12, 13]. Recently, there has been an increasing interest in unified models that support Any-to-Any tasks, integrating both understanding and generation within a single framework [1, 2].
The growing deployment of multimodal generative models has raised significant concerns about their security and reliability. Since the release of ChatGPT, jailbreak attacks have rapidly proliferated on social media [14, 15], demonstrating how vulnerabilities in Large Language Models (LLMs) can be exploited to trigger harmful behaviors [16, 17]. Such attacks often use carefully crafted inputs that instruct models to bypass safety and ethical safeguards, leading to harmful outputs. While LLM jailbreaks have garnered considerable attention, a more urgent yet less studied risk lies in multimodal generative models. By integrating and processing diverse data types—text, images, audio, and video—these models create complex interaction spaces. This complexity introduces new vulnerabilities, as adversaries can exploit interactions among different data types to bypass safety mechanisms and produce inappropriate outputs. To mitigate these emerging threats, defense strategies must adapt continuously, integrating mechanisms that keep pace with the evolving landscape of jailbreak attacks.
Figure 1: Illustrated examples of jailbreak attacks on multimodal generative models to induce harmful outputs across various modalities, including harmful text via the Jailbreak in Pieces [18], harmful images via the MMA-diffusion [19], harmful videos via the T2VSafetyBench [20] and harmful audio via the Voice Jailbreak [21].
Figure 2: Jailbreak attack and defense against multimodal generative models.
Existing reviews on jailbreak methods [22, 23, 24, 25] have primarily concentrated on content output within a specific modality, addressing tasks either Any-to-Text [26, 23] or Any-to-Image [24, 25]. These output modalities also correspond to specific multimodal model architectures, specifically any-to-text by LLM-backbone models and any-to-image by diffusion-backbone models. While these surveys have made valuable contributions, they lack a unified framework that spans a broad range of modalities (i.e., text, images, audio, and video) within different generative systems, as shown in Fig. 1.
To address this gap, our survey introduces the first unified framework which systematically summarizes jailbreak attacks and defense mechanisms across various input-output modalities and different generative structures. Specifically, we break down the lifecycle of multimodal jailbreak and abstract four discrete levels – input, encoder, generator, and output. This structured approach helps bridge the gaps between different models, each of which may have unique architecture but share common vulnerabilities within these four key levels. We outline the four general steps (as shown in Fig. 2) to devise jailbreak attack and defense techniques:
  1. 1.
    Input Level: Attackers and defenders operate solely on the input data. Attackers modify inputs to execute attacks, while defenders incorporate protective cues to enhance detection.
  2. 2.
    Encoder Level: With access to the encoder, attackers optimize adversarial inputs to inject malicious information into the encoding process, while defenders work to prevent harmful information from being encoded within the latent space.
  3. 3.
    Generator [^1] We refer to the entire generative model as the “Generator” without loss of generality. Level: With full access to the generative models, attackers leverage inference information, such as activations and gradients, and fine-tune models to increase adversarial effectiveness, while defenders use these techniques to strengthen model robustness.
  4. 4.
    Output Level: With the output from the generative model, attackers can iteratively refine adversarial inputs, while defenders can apply post-processing techniques to enhance detection.
Note that we encompass a broader range of input-output modality configurations, including text, image, audio, and video, alongside multiple types of multimodal generative models such as Any-to-Text, Any-to-Vision, and Any-to-Any models. Meanwhile, we conduct a comparative analysis of various evaluation datasets and metrics used for benchmarking, along with insightful observations and suggestions for future research directions. By highlighting the landscape of jailbreak attacks against multimodal generative models, our survey enhances the understanding of security challenges and provides direction for developing effective defenses. We aim to equip researchers, practitioners, and policymakers with valuable insights to safeguard foundation models against malicious exploitation. In summary, our key contributions are as follows:
  • Through a comprehensive review of existing attack methodologies (see TABLE III) and defense strategies (see TABLE IV), we abstract and summarize a general categorization for launching and defending jailbreak against multimodal generative models, comprising four distinct stages (see Fig. 2).
  • We present a comprehensive and systematic review of attack, defense, and evaluation strategies across various input-output modalities and different model structures.
  • We thoroughly discuss the limitations, challenges, and future directions for real-world applications, facilitating future research in this domain.
The remaining paper is organized as follows. We first provide a brief introduction to the preliminaries in Section II, which cover essential topics and concepts for the proper understanding of this work. In Section III and Section IV, we summarize existing approaches for jailbreak attacks and defense strategies, based on the stages of interaction with generative models respectively, including input-level, encoder-level, generator-level, and output-level. Section V introduces the commonly used datasets and evaluation metrics. Moreover, we provide discussions and future research opportunities in Section VI. Finally, we conclude this review in Section VII.
II PRELIMINARIES
In this section, we provide a concise introduction to multimodal generative models and jailbreak, aiming to enhance comprehension of our work.
II-A Multimodal Generative Models
Current multimodal generative models can be broadly classified into three distinct categories. The first category includes Any-to-Text (Any Modality to Text) Models, which integrate inputs from multiple modalities, encode them, and project into the word embedding space of the LLM for generating textual output[3, 27, 28, 4, 29, 5, 30]. The second category encompasses Any-to-Vision (Any Modality to Vision) Models, which encode inputs across different modalities as conditional information and leverage diffusion models to generate visual outputs [10, 31, 32, 33, 34, 35, 36, 37, 38]. Thirdly, Any-to-Any (Any Modality to Any Modality) Models perceive inputs and generate outputs in arbitrary combinations of text, image, video, and audio [39, 40, 41, 2]. We summarize the combinations of modalities regarding both inputs and outputs of all categories in TABLE I. Additionally, we provide a comprehensive analysis of their underlying architectures as follows:
TABLE I: Illustration of model short name and representative generative models used for jailbreak. For input/output modalities, I: Image, T: Text, V: Video, A: Audio, Any: Any of Text/Image/Video/Audio.
Short Name Modality Representative Model
Any-to-Text Models (LLM Backbone)
IT2T I + T → T LLaVA [3], MiniGPT4 [27], InstructBLIP [28]
VT2T V + T → T Video-LLaMA [4], Video-LLaVA [29]
AT2T A + T → T Audio Flamingo [5], AudioPaLM [30]
Any-to-Vision Models (Diffusion Backbone)
T2I T → I Stable Diffusion [10], Midjourney [31], DALLE [32]
IT2I I + T → I DreamBooth [33], InstructP2P [34]
T2V T → V Open-Sora [35], Stable Video Diffusion [36]
IT2V I + T → V VideoPoet [37], CogVideoX [38]
Any-to-Any Models (Unified Backbone)
IT2IT I + T → I + T Next-GPT [39], Chameleon [40]
TIV2TIV T + I + V → T + I + V EMU3 [2]
Any2Any Any → Any GPT-4o [41], Gemini Ultra [42]
$\bullet$ Any-to-Text Models. Typical models in this category consist of three primary components: an encoder, a pre-trained LLM, and a modality interface that connects them. The modality encoder functions akin to human sensory organs, transforming raw visual or audio data into compact representations. A common approach is to use pre-trained encoders that are already aligned with language data, as seen in CLIP models [43], which facilitate alignment with LLMs. The LLM, often chosen from established pre-trained models like LLaMA [44] and Vicuna [44], serves as the central reasoning unit. These models benefit from extensive pre-training on web corpora, allowing for rich knowledge representation and reasoning capabilities. To bridge the gap between modalities and language, a modality interface is introduced. This interface can either be a learnable projector that directly aligns the encoded modality features with the LLM’s input requirements or an expert model that translates non-textual data into language. Overall, Any-to-Text Models utilize a multi-module architecture to effectively integrate multimodal inputs and generate coherent textual outputs.
$\bullet$ Any-to-Vision Models. Diffusion models represent a major breakthrough in visual generation, surpassing traditional methods like Generative Adversarial Networks (GANs) [45]. They are extensively used in generative tasks. By operating as parameterized Markov chains, diffusion models transform random noise into complex images through sequential refinement. The reverse process is modeled as a normal distribution. A vision processor, often UNet, is trained to estimate the mean and variance.
For image generation, Ho et al.[46] introduce classifier-free guidance by training an additional text encoder as a condition interpreter, allowing for flexible conditioning with free-form text prompts. Beyond text, ControlNet[47] incorporates specific conditions like depth maps, normal maps, canny edges, poses, and sketches to enhance fidelity and controllability. Methods such as UniControlNet [48], UniControl [49], Composer [50], and T2I-Adapter [51] further improve generation by integrating diverse conditions and additional layers as condition interpreters.
Text-to-video generation models rely on three core components: condition interpreters, vision processors, and temporal handlers [52]. Condition interpreters translate the input text into visual elements, connecting the semantics of the text with the objects in the image and their dynamics in the video. Vision processors handle the visual content within each frame. Since a video is essentially a sequence of images, these models often use vision processing modules similar to those in image generation. Temporal handlers manage the progression between frames, learning the dynamics of visual content over time. This element is unique to text-to-video models, enabling the capture of motion and transitions between frames through various mechanisms.
$\bullet$ Any-to-Any Models. Typical models in this category can be grouped into two main approaches: the first integrates a foundational language model with an additional generator, such as a pre-trained diffusion model; the second employs a unified Transformer architecture that jointly manages both comprehension and generation. In the first approach, models like Next-GPT [39] utilize a pre-trained LLM to interpret multimodal inputs, by an independent diffusion model for image generation. The LLM functions as the core reasoning unit, while the diffusion model ensures the production of coherent visual outputs, thereby supporting complex multimodal tasks.
In contrast, the second approach embodies a fully integrated solution, where a single Transformer mode simultaneously processes both understanding and generation tasks across multiple modalities, thereby eliminating the dependence on diffusion or compositional methods. For instance, Chameleon [1] leverages interleaved multiple-modality tokens, allowing for joint reasoning over both modalities within a unified architecture. Recent research introduces Emu3 [2], a novel approach that relies solely on next-token prediction. EMU3 tokenizes text, images, and videos into a unified discrete space and utilizes a single transformer trained from scratch across a mixture of multimodal sequences.
TABLE II: Notations related to jailbreak against multimodal models. For input/output modalities, I: Image, T: Text, V: Video, A: Audio.
Symbol Description
$\mathcal{X}_{mal}$ Malicious input.
$\mathcal{X}_{adv}$ Adversarial input.
$\mathcal{Y}_{mal}$ Malicious output.
$\mathcal{Y}_{adv}$ Adversarial output.
$E_{M}$ Encoder, $M$ can be any of $\left\{T,I,V,A\right\}$.
$\mathcal{M}_{\theta}$ Generative Model.
$\mathcal{F}_{sc}$ Safety Checker.
$\mathcal{S}_{tox}$ Assessment of the input toxic scores.
$\mathcal{S}_{harm}$ Assessment of the output harmful scores.
$p_{\theta}\left(y|x\right)$ Probability of next token prediction.
$\epsilon_{\theta}\left(z_{t}|c\right)$ Predicted noise generated by Denoising Networks.
II-B Jailbreak Attack
$\bullet$ Problem Formulation. Jailbreak attacks against generative models $\mathcal{M}_{\theta}$ involve exploiting vulnerabilities to elicit unintended or harmful outputs $\mathcal{Y}_{mal}$. Directly using malicious prompts $\mathcal{X}_{mal}$ is often filtered by the safety mechanisms implemented in these models. Therefore, successful jailbreak attempts require carefully crafted adversarial inputs $\mathcal{X}_{adv}$ designed to bypass these built-in safety protocols.
$\max_{\mathcal{X}_{adv}}\;\mathbb{E}_{\boldsymbol{x}\sim\mathcal{X}_{adv}}[\mathcal{S}_{harm}(\mathcal{M}_{\theta}(x))],$ (1)
$\mathrm{s}.\mathrm{t}.\;\mathcal{S}_{tox}(x)<\epsilon,$ (2)
where $\mathcal{L}_{harm}$ quantifies the degree of harmfulness in the generated content $\mathcal{Y}_{mal}$, and $\mathcal{L}_{tox}$ assesses the manifest toxicity of the adversarial prompt. $\epsilon$ denotes the corresponding thresholds. The objective is to optimize the generation of highly harmful content while maintaining manifest toxicity below a certain threshold, ensuring it avoids detection and filtering.
$\bullet$ Multimodal Jailbreak. Research on jailbreak attacks against generative models has emerged from the field of LLMs [17, 53]. For instance, Zou et al. [54] proposes to append specific adversarial suffixes to malicious prompts, while AutoDAN [55] employs a sophisticated hierarchical genetic algorithm to autonomously generate subtle jailbreak prompts. As generative models evolve beyond text to include multimodal architectures such as vision-language and text-to-image models, the scope of jailbreak attacks has similarly expanded. In these settings, adversaries exploit not only textual prompts but also visual and multimodal embeddings, requiring more advanced defense mechanisms to ensure the ethical and secure use of these technologies. Continued research in this area is crucial to developing effective safeguards against adversarial manipulation while preserving the utility and accessibility of generative models across a wide range of applications.
II-C Jailbreak Defense
To address these vulnerabilities, significant efforts have focused on enhancing the safety alignment of LLMs by improving data quality during the Supervised Fine-tuning (SFT) phase [56] and aligning model behavior with human safety preferences during the Reinforcement Learning from Human Feedback (RLHF) stage [57]. Although LLMs have made significant progress in safety alignment, the introduction of additional modalities in multimodal generative models adds layers of complexity and exposes new vulnerabilities. For instance, while harmful text prompts alone may be blocked, malicious content embedded as typographic images can still bypass defenses and lead to the generation of unsafe content. This underscores the importance of robust jailbreak defense mechanisms tailored to the unique structure and applications of multimodal generative models.
III Jailbreak Attack
In this section, we focus on discussing different advanced jailbreak attacks against multimodal models. We categorize attack methods into black-box, gray-box, and white-box attacks (refer to TABLE III). As shown in Fig. 2, in a black-box setting where the model is inaccessible to the attacker, the attack is limited to surface-level interactions, focusing solely on the model’s input and/or output. Regarding gray-box and white-box attacks, we consider model-level attacks, including attacks at both the encoder and generator.
III-A Black-box Jailbreak
In black-box scenarios, attackers lack direct access to the internal architecture, parameter configurations, or gradient details of the target model, restricting their interactions to observing and manipulating only the model’s input-output behavior.
Figure 3: Illustration of black box jailbreak attacks against multimodal generative models at the input level where attackers focus on devising sophisticated jailbreak input templates.
TABLE III: Summary of jailbreak attack methods against multimodal models. Black means black-box attack, white means white-box attack, and gray means gray-box attack. For input/output modalities, I: Image, T: Text, V: Video, A: Audio.
Taxonomy
Method Venue Access Level Category Model Contribution
SASP [58] [arXiv’23] Black Input Prompt Engineering I + T → T Extract the system prompts from GPT-4V for jailbreaking MLLMs.
AutoJailbreak [59] [NACCLW’24] Black Input Prompt Engineering I + T → T Leverage LLM’s native prompt optimization to automate jailbreaks.
Arondight [60] [ACM MM’24] Black Input Prompt Engineering I + T → T Generate multimodal prompts via MLLMs/LLMs guided by reinforcement learning.
AdvWeb [61] [arXiv’24] Black Input Prompt Engineering I + T → T Optimize an adversarial prompter model to mislead MLLM-powered web agents.
AIAH [62] [arXiv’24] Black Input Prompt Engineering A + T → T Decompose harmful words into letters and conceal them in audio input.
Prompt Dilution [63] [NeurIPSW’22] Black Input Prompt Engineering T → I Dilute prompts by adding unrelated extra to degrade filter performance.
PGJ [64] [arXiv’24] Black Input Prompt Engineering T → I Use LLMs to find perceptually similar safe phrases to replace unsafe words.
SurrogatePrompt [65] [CCS’24] Black Input Prompt Engineering T → I Use LLMs to substitute high-risk sections within a suspect prompt.
APGP [66] [arXiv’24] Black Input Prompt Engineering T → I Use LLMs to refine prompts via score function for copyright infringement.
DACA [67] [arXiv’23] Black Input Prompt Engineering T → I Use LLMs as text transformation agents to generate adversarial prompts.
Atlas [68] [arXiv’24] Black Input Prompt Engineering T → I Develop three LLM agents with COT reasoning for T2I jailbreak.
UPAM [69] [ICML’24] Black Input Prompt Engineering T → I Optimize LLMs to generate natural adversarial prompts via two-stage learning.
BSPA [70] [arXiv’24] Black Input Prompt Engineering T → I Optimize text retriever to identify sensitive words.
Figstep [71] [arXiv’23] Black Input Image Engineering I + T → T Transform malicious text prompts into image form overlaid on white backgrounds.
MM-SafetyBench [72] [ECCV’24] Black Input Image Engineering I + T → T Use Stable Diffusion to generate images based on the extracted keywords.
Logic Jailbreak [73] [arXiv’24] Black Input Image Engineering I + T → T Design flowchart images to access MLLM’ logical reasoning abilities.
Visual-RolePlay [74] [arXiv’24] Black Input Role Play I + T → T Act as high-risk roles in image inputs to generate harmful content.
Voice Jailbreak [21] [arXiv’24] Black Input Role Play A → A Use fictional storytelling elements in voice prompts to jailbreak GPT-4o.
DiffZOO [75] [arXiv’24] Black Output Estimation-based T → I Estimate gradients via zero-order optimization for crafting adversarial prmpts.
SneakyPrompt [76] [S&P’24] Black Output Search-based T → I Use reinforcement learning to guide the search process for refining prompts.
Jailbreak in Pieces [18] [ICLR’24] Gray Encoder Gradient-based I + T → T Combine adversarial images within benign prompts to disrupt safety alignment.
Redteaming Attack [77] [ECCV’24] White Encoder Gradient-based I + T → T Mislead VLLM outputs by attacking on the vision encoder.
MMA-Diffusion [19] [CVPR’24] White Encoder Gradient-based T → I Generate adversarial prompts via gradient optimization and word regularization.
JPA [78] [arXiv’24] White Encoder Gradient-based T → I Optimize prefix prompts in latent space to align malicious concepts.
Ring-A-Bell [79] [ICLR’24] Gray Encoder Search-based T → I Extract holistic concepts within latent space and employ genetic search.
RT-Attack [80] [arXiv’24] Gray Encoder Search-based T → I Two-stage random search to optimize prompt-wise and image-wise similarity.
Image Hijacks [81] [arXiv’24] White Generator Gradient-based I + T → T Optimize image hijacks using behavior matching for multi-type attacks.
VisualAdv [82] [AAAI’24] White Generator Gradient-based I + T → T Optimize a single adversarial example on few-shot harmful corpus.
ImgJP [83] [arXiv’24] White Generator Gradient-based I + T → T Generate adversarial images based on maximum likelihood algorithm.
ImgJS [84] [arXiv’24] White Generator Gradient-based I + T → T Convert visual adversarial vectors to text space, merged with harmful queries.
Adv Aligned [85] [NeurIPS’23] White Generator Gradient-based I + T → T Optimize adversarial images to increase the probability of harmful response.
HADES [86] [ECCV’24] White Generator Gradient-based I + T → T Composite images integrating malicious text, harmful images, adversarial noise.
Agent Smith [87] [ICML’24] White Generator Gradient-based I + T → T Inject adversarial images in multi-agent memory to elicit harmful responses.
UMK [88] [arXiv’24] White Generator Gradient-based I + T → T Optimize adversarial prefixed across image-text modalities simultaneously.
BAP [89] [arXiv’24] White Generator Gradient-based I + T → T Combine query-agnostic image perturbing and intent-specific text optimization.
P4D [90] [ICML’24] White Generator Gradient-based T → I Optimize latent noise predictions to find the safety-evasive prompts.
UnlearnDiffAtk [91] [ECCV’24] White Generator Gradient-based T → I Optimize adversarial prompts by maximizing diffusion model likelihood.
VA3 [92] [CVPR’24] White Generator Gradient-based T → I Iterative prompt optimization using multi-armed bandit for copyright evasion.
MMA-Diffusion [19] [CVPR’24] White Generator Gradient-based I + T → I Optimize gradients of loss items that exceed the safety checker’s thresholds.
A.1 Input-level Attack
As shown in Fig. 3, attackers are compelled to develop more sophisticated input templates across prompt engineering, image engineering, and role-ploy techniques. These techniques can bypass the model’s safeguards, making the models more susceptible to executing prohibited instructions.
Prompt Engineering. Prompt engineering involves strategically modifying specific words or rewriting complete prompts. We broadly divide this approach into three types based on prompt sources: manual construction, generation using off-the-shelf LLM models, and generation using trained proxy models.
$\bullet$ Image + Text → Text (IT2T) Models. SASP [58] employs carefully crafted dialogues to extract confidential system prompts, facilitating the discovery of potential jailbreak prompts. Furthermore, these attack prompts are incorporated with manual modifications to enhance the attack success rate. Compared with manual construction, AutoJailbreak [59] leverages LLMs as red-team tools to automatically refine attack prompts. This framework integrates weak-to-strong prompt optimization strategies, a suffix-based attack enhancement technique, and an efficient search mechanism. Instead of relying on off-the-shelf LLMs, other methods involve optimizing proxy models for prompt refinement. Arondight [60] automates multimodal jailbreak attacks by generating textual prompts using a trained LLM, supplemented with toxic images crafted by the proprietary GPT-4V. The red team LLM, guided by a reinforcement learning agent, is optimized by incorporating entropy bonuses and novelty reward metrics to enhance prompt diversity. Similarly, AdvWeb [61] proposes the first black-box framework specifically designed to mislead MLLM-powered web agents. This framework optimizes a generative model to produce adversarial prompts, which are injected into web pages to execute targeted malicious actions.
$\bullet$ Audio + Text → Text (AT2T) Models. AIAH [62] introduces a speech-specific jailbreaking technique, that breaks down harmful words into individual letters to obscure their presence in the audio input. The model is then instructed to concatenate these letters back into complete words, thereby reconstructing the original harmful question within the jailbreak prompt.
$\bullet$ Text → Image (T2I) Models. Earlier work [63] proposes a “prompt dilution” strategy to degrade filter performance by adding an unrelated extra to prompts. Through reverse engineering, it is shown that the filter specifically targets sexual content while ignoring other problematic themes like violence.
Advances in off-the-shelf LLMs have also reduced the barrier to producing diverse jailbreak prompts on T2I models. By utilizing the concept of “substitution”, both PGJ [64] and SurrogatePrompt [65] employ LLMs, such as ChatGPT, to replace unsafe content with benign phrases. These approaches typically introduce perceptual confusion by preserving visual similarity while altering the textual semantics (e.g., replacing “blood” with “watermelon juice”). Unlike replacement-based methods, the automated Prompt Generation Pipeline (APGP) [66] utilizes LLMs to refine complete jailbreaking prompts by optimizing a self-generated score and imposing penalties for certain keywords, thereby effectively assessing copyright infringement risks in T2I systems. Furthermore, agent-based methods can unleash the potential of LLMs in prompt engineering. Divide-and-Conquer Attack (DACA) [67] leverages GPT models as text transformation agents to create adversarial prompts. They design attack helper prompts that guide LLMs to break down an unethical drawing intent into multiple benign descriptions of individual image elements. Building on a collaborative multi-agent framework, Atlas [68] employs two specialized agents: the mutation agent and the selection agent. Each agent is equipped with four modules: a MLLM/LLM brain, planning, memory, and tool usage. These two agents collaborate in an iterative process to determine jailbreak prompt-triggered responses and generate new candidate jailbreak prompts.
In a similar vein, training proxy models can support the development of specialized red-team tools for jailbreak-prompt generation in T2I models. The Unified Prompt Attack Model (UPAM)[69] aims to optimize LoRA adapters within LLMs to generate natural adversarial prompts by utilizing a two-stage learning framework. In the first stage, sphere-probing learning optimizes prompts to bypass defenses, while the second stage employs semantic-enhancing learning to refine the generated images to align with harmful targets. Additionally, BPSA[70] optimizes a text retriever to identify sensitive words associated with the input vector and leverages an LLM API to facilitate the generation of stealthy attack prompts.
Image Engineering. Image engineering methods primarily embed toxic content into carefully crafted images by text typography and diffusion-based generation. Due to the insufficient security training in the visual modules of current generative models [71], it is relatively more vulnerable to jailbreak attacks.
$\bullet$ Image + Text → Text (IT2T) Models. Figstep [71] proposes to transform malicious text prompts into image-based typographic forms. This approach converts toxic questions into textual statements and then overlays these statements onto a white background image, along with indices as cues for response formatting. These images are subsequently paired with benign text prompts to execute the jailbreak attack. Some other methods have explored the potential of using diffusion models to generate targeted jailbreak images. MM-SafetyBench [72] employs a dual-approach strategy to generate images from malicious text prompts. The first approach utilizes stable diffusion [10] to generate images based on the extracted keywords from the malicious prompts. The second approach converts the keywords or entities into visual typographic forms. Additionally, the Logical jailbreak method [73] designs flowchart images with semantic meaning to evaluate the logical reasoning and visual imagination capabilities of MLLMs.
Role Play. Role-play methods often construct targeted scenarios that induce the model to take on a role aligned with harmful behaviors, leading to successful jailbreaking attempts. This method leverages the model’s capacity for complex interactions and role simulation, often through visually-driven or story-driven roles.
$\bullet$ Image + Text → Text (IT2T) Models. Visual-RolePlay [74] introduces a role-playing framework that generates images depicting high-risk roles, with descriptive text positioned at the top and malicious prompts at the bottom. After generating these role-specific images, the model is instructed to play high-risk roles in image inputs to produce harmful content.
$\bullet$ Audio → Audio (A2A) Models. VoiceJailbreak [21] introduces an innovative voice-based jailbreak attack on GPT-4o by exploiting fictional storytelling techniques. This method constructs prompts around key elements of fictional writing – such as setting, character, and plot. By engaging the model as a reader of fictional narratives, the adversary rephrases forbidden questions as assertive statements within fictional contexts, achieving a higher attack success rate through multi-step voice inputs.
Figure 4: Illustration of black box jailbreak attacks against multimodal generative models at the output level where attackers focus on querying responses to multiple input variants to obtain satisfactory jailbreak outputs.
A.2 Output-level Attack
In Fig. 4, attackers focus on querying outputs across multiple input variants, systematically analyzing the model’s responses to diverse inputs. Driven by specific adversarial goals, attackers employ estimation-based and search-based attack techniques to iteratively refine these input variants.
Estimation-based Attack. Estimation-based methods typically involve approximating gradients through zeroth-order estimation techniques, where attackers rely on black-box queries to estimate the model’s gradient by observing output variations in response to small input perturbations.
$\bullet$ Text → Image (T2I) Models. DiffZOO[75] addresses the challenges of gradient unavailability and discrete prompt optimization by employing zeroth order optimization to estimate gradients. During the optimization process, DiffZOO identifies critical tokens within prompts that significantly impact model behavior and selectively replaces them to enhance the attack’s effectiveness.
Search-based Attack. Search-based methods are employed to efficiently navigate the candidate input space, refining input variations to align with specified adversarial objectives.
$\bullet$ Text → Image (T2I) Models. SneakyPrompt [76] is an automated attack framework that leverages reinforcement learning to assist in replacing tokens within the search space. Reinforcement learning is used to guide the search process, optimizing adversarial jailbreak prompts by iteratively refining them based on reward signals. The framework’s reward model is grounded in cosine similarity between adversarial and malicious images in the CLIP embedding space, effectively steering the optimization toward generating unsafe content.
III-B White-box and Gray-box Jailbreak
In a white-box attack scenario, attackers have access to the internal architecture of the target model, including encoder and generator modules. This level of access allows attackers to leverage gradient information and the model’s intermediate representations, facilitating precise adversarial modifications. In certain scenarios, attackers possess partial knowledge of the model, such as its output distribution or structural details. This type of attack is referred to as a gray-box attack.
Figure 5: Illustration of white-box jailbreak attacks against multimodal generative models at the encoder level where attackers can directly exploit the vulnerabilities within the encoder architecture.
B.1 Encoder-level Attack
For encoder-level attacks, attackers are restricted to accessing only the encoders to provoke harmful responses. In this case, attackers typically seek to maximize cosine similarity within the latent space, ensuring the adversarial input retains similar semantics to the target malicious content while still being classified as safe. This adversarial objective is formalized as:
$\underset{\mathcal{X}_{adv}}{\mathrm{arg}\max}\,\,Cos\left(E_{M}\left(\mathcal{X}_{adv}\right),E_{M}\left(\mathcal{X}_{mal}\right)\right).$ (3)
This objective can be solved using gradient-based or search-based optimization techniques in Fig. 5.
Gradient-based Attack. Gradient-based attacks operate by extracting gradient information from the model to iteratively modify the input, moving the input closer to achieve the adversarial objective.
$\bullet$ Image + Text → Text (IT2T) Models. Jailbreak in Pieces [18] and Redteaming Attack [77] both propose to construct adversarial images paired with textual prompts to disrupt model alignment. These adversarial images are optimized through gradient-based attacks within the latent space of the encoder, effectively mapping adversarial embeddings near regions associated with harmful triggers. Notably, Jailbreak in Pieces [18] targets the CLIP [43] as the surrogate encoder, which shares the same model architecture as the built-in encoder in MLLMs.
$\bullet$ Text → Image (T2I) Models. For T2I models, both MMA-Diffusion [19] and JPA [78] employ gradient-driven techniques to iteratively optimize adversarial prompts, aligning their embeddings closely with those of the target harmful content. Additionally, MMA-Diffusion [19] applies word regularization methods to remove any sensitive terms, while JPA [93] merges inappropriate concepts into the target content embedding to counter concept removal methods.
Search-based Attack (Gray-box). In certain scenarios, attackers may lack full white-box gradient access but retain access to model output distributions, such as encoder-generated embeddings. These embeddings can be exploited to construct adversarial objectives, with search-based methods used for optimization to avoid explicit gradient requirements.
$\bullet$ Text → Image (T2I) Models. Ring-A-Bell [79] iteratively explores the latent space, adjusting adversarial inputs to align their embeddings with problematic prompts. To achieve this, the method employs a genetic algorithm [94] to optimize input modifications, enabling a guided exploration of semantic similarity. To expand the search space, RT-Attack [80] aims to find the adversarial prompt within a vast vocabulary codebook. This method utilizes a naive random search strategy guided by a two-stage adversarial objective – first optimizing semantic similarity between the adversarial and target malicious prompts, then maximizing the similarity between images generated by two prompts.
Figure 6: Illustration of white-box/gray-box jailbreak attacks against multimodal generative models at the generator level where attackers have full access to the entire model architecture.
B.4 Generator-level Attack
For generator-level attacks in Fig. 6, attackers have unrestricted access to the generative model’s architecture and checkpoint, enabling attackers to conduct thorough investigations and manipulations, thus enabling sophisticated attacks.
Gradient-based Attack. Here gradient-based attacks are driven by defined adversarial optimization objectives by using gradients derived from the full generative models to refine adversarial inputs. Based on the source of supervisory signals embedded within the optimization objective, gradient-based attacks can be categorized into three types: target response-based, auxiliary generative model-based, and auxiliary classifier-based.
$\bullet$ Image + Text → Text (IT2T) Models. Most existing methods [81, 82, 83, 84, 85, 86, 87, 88, 89] operate gradient-based attacks against MLLMs typically are driven by the target responses, by maximizing the likelihood loss to achieve targeted modifications in the output probability distribution. This is formally expressed as:
$\mathop{\mathrm{arg}\min}\limits_{\mathcal{X}_{adv}}\;-\sum_{i=1}^{m}{\log\left(p_{\theta}\left(\mathcal{Y}_{i}|\mathcal{X}_{adv}\right)\right)},$ (4)
where $\left\{\mathcal{Y}_{i}\right\}_{i=1}^{m}$ represents the target output sequence, $m$ is the sequence length and $p_{\theta}\left(\cdot\right)$ represents the probability of next token prediction.
Specifically, Image Hijacks [81] crafts prompt-matching adversarial images to match malicious behaviors to arbitrary text prompts. VisualAdv [82] reveals that a single visual adversarial example generated using a “few-shot” toxic corpus can universally jailbreak aligned models. In addition to adversarial noise as a visual component, HADES [86] introduces two additional elements: images created from harmful text typography and malicious images generated using stable diffusion [10]. Moreover, Agent Smith [87] addresses the issue of “infectious jailbreak” within a multi-agent environment. By introducing adversarial images into the memory of randomly selected agents, this method can induce the agents to generate harmful responses. In contrast to methods that primarily target the visual modality, UMK[88] and Bi-Modal Adversarial Prompt Attack (BAP) [89] conduct simultaneous adversarial attacks across both image and text modalities.
$\bullet$ Text → Image (T2I) Models. Given that victim T2I models $\mathcal{M}_{\theta_{V}}$ possess safety mechanisms that can remove the sensitive concept, existing methods [90, 91, 92] often perform gradient-based adversarial attacks aiming to minimize the noise prediction loss. For instance, Prompting4Debugging (P4D) [90] generates an image with an inappropriate concept by aligning noise predictions between victim T2I models $\mathcal{M}_{\theta_{V}}$ and unconstrained T2I models $\mathcal{M}_{\theta_{U}}$, which can be formulated as:
$\mathop{\mathrm{arg}\min}\limits_{\mathcal{X}_{adv}}\,\,\left\|\epsilon_{\theta_{U}}\left(z_{t}|\mathcal{X}_{mal}\right)-\epsilon_{\theta_{V}}\left(z_{t}|\mathcal{X}_{adv}\right)\right\|_{2}^{2}.$ (5)
Since two separate diffusion processes will increase the computational burden, UnlearnDiffAtk [91] prepare a target image $I_{tgt}$ containing sensitive content as optimized guidance to minimize discrepancies in noise predictions by:
$\mathop{\mathrm{arg}\min}\limits_{\mathcal{X}_{adv}}\,\,\left\|\epsilon-\epsilon_{\theta_{V}}\left(I_{tgt,t}|\mathcal{X}_{adv}\right)\right\|_{2}^{2}.$ (6)
Similarly, by simultaneously introducing target images and auxiliary models, VA3 [92] explores adversarial attacks in an online scenario where attackers iteratively generate prompts to increase the likelihood of copyright infringement. To enhance attack success, the proposed framework utilizes an amplification strategy and adversarial prompt optimization, leveraging a multi-armed bandit approach for prompt selection.
$\bullet$ Text + Image → Image (IT2I) Models. Apart from using auxiliary generative models or target responses as guidance, MMA-Diffusion [19] employs an external image classifier to produce classification supervision. Specifically, MMA-Diffusion [19] launches adversarial attacks on image editing tasks and dynamically optimizes the gradients of loss items that exceed the victim safety checker’s $\mathcal{F}_{sc}$ thresholds $T$, the method aims to minimally alter image features while bypassing the safety mechanism. This adversarial objective is formalized as:
$\mathop{\mathrm{arg}\min}\limits_{\mathcal{X}_{adv}}\,\,1_{\left\{Cos\left(f_{adv},f_{mal}\right)>T\right\}}Cos\left(f_{adv},f_{mal}\right),$ (7)
$f_{adv}=\mathcal{F}_{sc}\left(\mathcal{Y}_{adv}\right),\,\,f_{mal}=\mathcal{F}_{sc}\left(\mathcal{Y}_{mal}\right).$ (8)
IV Jailbreak Defense
In this section, we introduce current efforts made in the jailbreak defense of multimodal generative models, which includes two lines of work: Discriminative defense and Transformative defense (as shown in TABLE IV). In a discriminative setting, the defense is constrained to classification tasks for assigning binary labels. In contrast, transformative defense extends beyond classification to influencing the model’s generative process. Defenders aim to produce appropriate and safe responses in the presence of malicious or adversarial inputs.
IV-A Discriminative Defense
Discriminative defenses focus on identifying and analyzing varying classified cues, such as statistical information at the input level, embeddings at the encoder level, activations at the generator level, and response discrepancies at the output level. Since these defenses operate independently of the generation pipeline, they preserve the model’s structural integrity and ensure that its generative capabilities remain unaffected.
Figure 7: Illustration of discriminative jailbreak defense at four different levels.
TABLE IV: Summary of jailbreak defense methods against multimodal models. For input/output modalities, I: Image, T: Text, V: Video, A: Audio.
Taxonomy
Method Venue Function Level Category Model Contribution
Intra-Entropy Gap [95] [arXiv’24] Discriminative Input Statistics-based I + T → T Measure the entropy gap to detect visual anomalies indicative.
Blacklist [31, 96, 32] [Arxiv’23] Discriminative Input Statistics-based T → I Restrict the predefined set of words or phrases in the input prompts.
CIDER [97] [arXiv’24] Discriminative Encoder Embedding-based I + T → T Identify similarity differences between harmful text and denoising adversarial images.
Latent Guard [98] [ECCV’24] Discriminative Encoder Embedding-based T → I Learn a latent space for detecting harmful concepts in the input text embeddings.
GuardT2I [99] [NeurIPS’24] Discriminative Encoder Embedding-based T → I First use generative models to interpret intentions behind adversarial latent embeddings.
NEARSIDE [100] [arXiv’24] Discriminative Generator Embedding-based I + T → T Obtain attack direction based on benign and adversarial embeddings for classification.
JailGuard [101] [Arxiv’23] Discriminative Output Query-based I + T → T Mutate inputs and leverage response discrepancies to detect attacks.
Espresso [102] [arXiv’24] Discriminative Output Detector-based T → I Present a content detector based on Contrastive Language-Image Pre-training.
AdaShield [103] [Arxiv’24] Transformative Input Prefix-based I + T → T Devise general defense prompts and prepend them to model inputs.
BlueSuffix [104] [Arxiv’24] Transformative Input Prefix-based I + T → T Fine-tune a generator through reinforcement learning to generate defensive suffix.
UNIGUARD [105] [Arxiv’24] Transformative Input Prefix-based I + T → T Search for additive visual noise and textual suffix to purify adversarial inputs.
POSI [106] [arXiv’24] Transformative Input Refiner-based T → I Fine-tune a LLM as an optimizer which transforms toxic prompts into clean ones.
Sim-CLIP [107, 108] [arXiv’24] Transformative Encoder Tuning-based I + T → T Fine-tune vision encoder through adversarial training.
AdvUnlearn [109] [NeurIPS’24] Transformative Encoder Tuning-based T → I Propose bi-level optimization-based integration scheme to fine-tune text encoder.
Safe-CLIP [110] [ECCV’24] Transformative Encoder Tuning-based T → I Fine-tune the CLIP model on synthetic data obtained from a LLM.
Self-discovery [111] [CVPR’24] Transformative Encoder Guidance-based T → I Discover vectors representing desired concepts to guide responsible generation.
VLGuard [112] [ICML’24] Transformative Generator Tuning-based I + T → T Build the first safety fine-tuning datasets for fine-tuning VLMs.
SafeVLM [113] [arXiv’24] Transformative Generator Tuning-based I + T → T Fine-tune additional safety modules and VLMs progressively in two stages.
Bathe [114] [arXiv’24] Transformative Generator Tuning-based I + T → T Fine-tune a trigger to connect harmful instructions with rejection responses.
Unlearn [115] [arXiv’24] Transformative Generator Tuning-based I + T → T Fine-tune VLMs via the unlearning method applied within the textual domain.
Receler [116] [ECCV’24] Transformative Generator Tuning-based T → I Propose concept-localized regularization and adversarial prompt learning.
EraseDiff [117] [arXiv’24] Transformative Generator Tuning-based T → I Fine-tune the model with remaining data and forgetting data.
Forget-Me-Not [118] [CVPR’24] Transformative Generator Tuning-based T → I Fine-tune the UNet to minimize the attention maps corresponding to the target concepts.
SalUn [119] [ICLR’24] Transformative Generator Tuning-based T → I Design weight saliency which apply machine unlearning to specific influential weights.
UCE [120] [WACV’24] Transformative Generator Tuning-based T → I Use a closed-form solution to modify the model’s attention weights.
MACE [121] [CVPR’24] Transformative Generator Tuning-based T → I Combine closed-form cross-attention refinement with LoRA fine-tuning.
ESD [122] [ICCV’23] Transformative Generator Tuning-based T → I Fine-tune the model by using conditioned and unconditioned scores.
SDD [123] [ICMLW’23] Transformative Generator Tuning-based T → I Apply self-distillation to fine-tune the diffusion model.
SafeGen [124] [CCS’24] Transformative Generator Tuning-based T → I Regulate the vision-only self-attention layers to remove concepts.
SFD [125] [arXiv’24] Transformative Generator Tuning-based T → I Align the conditional scores of “unsafe” classes or concepts with those of “safe” ones.
UC [126] [arXiv’24] Transformative Generator Tuning-based T → I Propose domain correction framework using adversarial training.
DUO [127] [ICMLW’24] Transformative Generator Tuning-based T → I Use preference optimization to fine tune the model.
RECE [128] [ECCV’24] Transformative Generator Tuning-based T → I Use closed-form parameter editing combined with adversarial learning schemes.
Dark Miner [129] [arXiv’24] Transformative Generator Tuning-based T → I Remove unsafe concept via mining, verifying, and circumventing.
ShieldDiff [130] [arXiv’24] Transformative Generator Tuning-based T → I Design a content-safe reward function to fine-tune the model via reinforcement learning.
InferAligner [131] [arXiv’24] Transformative Generator Guidance-based I + T → T Extract safety steering vectors to modify the activations of the victim model.
EIUP [132] [arXiv’24] Transformative Generator Guidance-based T → I Identiey and re-weight non-compliant features through attention mechanism.
SLD [93] [CVPR’23] Transformative Generator Guidance-based T → I Guide generation in the opposite direction of ”unsafe” prompt via classifier-free guidance.
Geom-Erasing [133] [ECCV’24] Transformative Generator Guidance-based T → I Extracts geometric information of undesire concepts that serve as negative prompts.
SAFREE [134] [arXiv’24] Transformative Generator Guidance-based T → V Dynamically adjust the denoising steps and design re-attention mechanisms.
P-ESD [135] [arXiv’24] Transformative Generator Pruning-based T → I Use model pruning to remove critical parameters linked to the unsafe concepts.
ConceptPrune [136] [arXiv’24] Transformative Generator Pruning-based T → I Identify neurons responsible for undesirable concepts.
MLLM-Protector [137] [arXiv’24] Transformative Output Refiner-based I + T → T Use detectors and detoxifiers to convert harmful responses to benign.
LMIVS [138] [WACV’24] Transformative Output Refiner-based T → I Use detectors and generators to detect and manipulate visual immorality.
A.1 Input-level Defense
Input analysis typically involves examining statistical features of the input, including detecting sensitive keywords, and assessing perplexity levels.
Statistics-based Defense. Statistical-based methods typically rely on established statistical metrics and rule-based criteria rather than complex models, making them easier to implement and interpret in various practical scenarios.
$\bullet$ Image + Text → Text (IT2T) Models. Intra-Entropy Gap [95] targets the detection of non-stealthy manipulations. This method analyzes entropy and perplexity-based differences between non-overlapping regions in image or text data to detect inconsistencies indicative of an attack.
$\bullet$ Text → Image (T2I) Models. In commercial T2I products such as Midjourney [31], Leonardo.ai [96] and DALL-E [32], blacklist methods are a commonly employed strategy for ensuring content safety and compliance with platform guidelines. These methods involve creating a predefined list of restricted or prohibited words and phrases that the T2I model must filter out or modify in its inputs.
A.2 Encoder-level Defense
Encoder-level defense focuses on the examination of the latent space within the encoder of generative models to assess the semantic tendencies of latent representations.
Embedding-based Defense. Without relying solely on explicit keywords or rules, embedding-based defenses assess the semantics and similarities between embeddings.
$\bullet$ Image + Text → Text (IT2T) Models. CIDER [97] performs iterative denoising of the image input and calculates the cross-modal similarity between the text and denoised image embeddings. If the difference in similarity surpasses a predefined threshold, the input is classified as adversarial, and the model responds by rejecting the request.
$\bullet$ Text → Image (T2I) Models. Latent Guard [98] detects unsafe T2I prompts by mapping the latent representations of blacklisted concepts and corresponding prompts within a shared latent space. The approach begins by utilizing a pre-trained text encoder to extract embeddings, followed by training an Embedding Mapping Layer that emphasizes relevant tokens through cross-attention mechanisms. Similarly, GuardT2I [99] optimizes a generative framework to convert latent embeddings into natural language, accurately capturing the user’s intent. By revealing the true intent of input prompts, GuardT2I then applies both blacklist and Sentence Similarity Checker to detect malicious prompts.
A.3 Generator-level Defense
Generator-level defense involves distilling internal embeddings or analyzing the probability distributions to identify anomalous patterns indicative of harmful content.
Embedding-based Defense. These defenses leverage embedding values from the hidden states of the generative model to monitor suspicious activity.
$\bullet$ Image + Text → Text (IT2T) Models. NEARSIDE [100] introduces a method to detect adversarial inputs by identifying the attack direction within the hidden space of VLMs. This attack direction is derived by calculating the average embedding difference between benign and adversarial inputs. An input is then classified as adversarial if the projection of its embedding onto the obtained attack direction exceeds a defined threshold.
A.4 Output-level Defense
Output-level Defenses typically employ iterative querying and specialized toxic detectors to assess the final outputs of the model.
Query-based Defense. Query-based methods involve repeatedly querying the model with perturbed inputs and analyzing the variance in the resulting outputs.
$\bullet$ Image + Text → Text (IT2T) Models. JailGuard [101] leverages input mutation and response divergence to identify attacks. The framework implements 18 mutators to mutate untrusted inputs, generating multiple variants, and distinguishing between adversarial and benign inputs based on the semantic divergence of their response patterns.
Detector-based Defense. To enhance detection accuracy, an alternative approach is to train a dedicated toxicity classifier.
$\bullet$ Text → Image (T2I) Models. Espresso [102] employs a fine-tuned CLIP classifier to filter unacceptable content in generated images. The fine-tuning process separates the embeddings of acceptable and unacceptable concepts, optimizing the cosine similarity for both while preserving contextual information.
Figure 8: Illustration of transformative jailbreak defense at four different levels. By influencing the generation process when encountering attacks, generative models produce safe response examples including text via AdaShield [103] and images via Forget-Me-Not [118].
IV-B Transformative Defense
Despite the improved accuracy in discriminative detection methods, these approaches often incorrectly flag benign inputs as harmful. Consequently, there is a growing research focus on transformative defenses that can operate at four levels to influence the model’s generation process, ensuring benign responses even in the presence of adversarial or malicious prompts in Fig. 8.
B.1 Input-level Defense
Defenders can enhance input prompts by embedding defense cues at the input level, which act as safeguards to steer the model towards generating benign outputs when confronted with malicious prompts.
Prefix-based Defense. Prefix-based Defense entails appending carefully crafted prefixes to input prompts as a preventative measure, guiding the model toward generating benign content.
$\bullet$ Image + Text → Text (IT2T) Models. AdaShield [103] optimizes a prompt generator to construct a diverse pool of defense prompts tailored to various scenarios. During inference, this approach selects the most suitable defense prompt from the pool based on semantic similarity, appending it as a prefix to model inputs to guard against jailbreak attacks. Moreover, BlueSuffix [103] comprises three key components: a suffix generator fine-tuned through reinforcement learning, a textual purifier, and a visual purifier. Specifically, the suffix generator produces a defensive suffix, which is appended to the purified text input generated by the textual purifier. These modified text inputs, combined with the purified images from the visual purifier, are then used as defense inputs for the target VLMs. Unlike approaches that use proxy models to generate prefixes, UNIGUARD [105] applies gradient-based optimization to construct specialized safety guardrails for each modality. These guardrails including visual noises and textual suffixes, are optimized to minimize the likelihood of generating harmful responses in a toxic corpus.
Refiner-based Defense. Refiner-based Defense typically requires training a dedicated generative model, specifically designed to transform harmful content into safe ones.
$\bullet$ Text → Image (T2I) Models. POSI [106] uses Proximal Policy Optimization to fine-tune a language model as an optimizer, transforming toxic prompts into clean ones. It introduces a novel reward function that assesses both the toxicity level and the textual alignment of the generated images.
B.2 Encoder-level Defense
For transformative defense, encoder-level methods focus on modifying the encoder’s behavior and can be categorized into two types: fine-tuning during the training phase and guidance implementation during the inference phase.
Tuning-based Defense. Fine-tuning encoders aims to achieve safety alignment within the model’s latent space, preventing the propagation of harmful content through the downstream generative process.
$\bullet$ Image + Text → Text (IT2T) Models. Sim-CLIP series methods [107, 108] introduce an unsupervised adversarial fine-tuning approach for the CLIP vision encoder to enhance robustness against adversarial attacks. The approach utilizes a siamese architecture, optimizing adversarial perturbations by maximizing cosine similarity between clean and perturbed image representations.
$\bullet$ Text → Image (T2I) Models. AdvUnlearn [109] proposes a utility-retaining regularization to optimize the trade-off between concept erasure robustness and model utility. It identifies the text encoder as more effective than UNet for robust unlearning and demonstrates its effectiveness in erasing nudity, objects, and style concepts across various diffusion models. Safe-CLIP [110] fine-tunes the CLIP model on synthetic data obtained from a LLM. It enhances safety in cross-modal tasks, including image-to-text retrieval, text-to-image generation, and image generation.
Guidance-based Defense. Guidance-based methods at the encoder level aim to embed specific information into the latent embeddings associated with potentially harmful content, steering the model’s behavior during content generation.
$\bullet$ Text → Image (T2I) Models. Self-discovery[111] proposes to operate on prompts implicitly. This approach discovers vectors representing desired concepts without requiring labeled data or external models. These vectors can be added to the embeddings of the condition prompt to guide responsible generation, promoting fairness and safety.
B.3 Generator-level Defense
Generative-level defense involves editing the model during training and incorporating guidance or pruning mechanisms during inference, thereby directly influencing the model’s internal generation process.
Tuning-based Defense. These approaches involve modifying the model’s behavior during the training time through adjustments to parameters or optimizing objectives, such as revising harmful prompts, erasing unsafe concepts, or strategically employing safety modules.
$\bullet$ Image + Text → Text (IT2T) Models. A group of meth- ods [112, 113, 114, 115] primarily fine-tune models with supervised instruction datasets to ensure the model’s output probability distribution aligns more closely with desired safe outputs. This is formally expressed as:
$\mathop{\mathrm{arg}\min}\limits_{\theta}\;-\sum_{i=1}^{m}{\log\left(p_{\theta}\left(\mathcal{Y}_{i}|\mathcal{X}_{harm}\right)\right)}.$ (9)
Specifically, VL-Guard [112] presents a vision-language safety dataset covering both harmful and safe images paired with relevant instructions. This dataset is used to evaluate and fine-tune models with two types of methods: post-hoc fine-tuning and mixed fine-tuning. Apart from solely training MLLMs, SafeVLM [113] introduces additional safety modules including a safety projector, safety tokens, and a safety head. Through a two-stage training process, this approach progressively trains safety modules and fine-tunes the language model. Similarly, BaThe [114] employs a series of trainable soft embeddings as a “wedge” within MLLMs. This approach maps harmful instructions to rejection responses during training, effectively transforming the model into a defense mechanism against jailbreak attacks.
Based on supervised fine-tuning, unlearning training involves removing specific harmful or unwanted information from a model through targeted retraining. Unlearn [115] proposes an unlearning approach by introducing loss terms that reduce harmful outputs while maintaining the model’s utility for benign inputs. This method focuses on three objectives: reducing the likelihood of generating harmful responses, encouraging helpful outputs in response to harmful inputs, and preserving the model’s original performance on benign inputs.
$\bullet$ Text → Image (T2I) Models. Fine-tuning pre-trained T2I models to erasure unsafe concepts has received wide attention. To erase the concept, the model is trained to predict the negatively guided noises that move the model’s prediction away from the erased concept.
$\mathop{\mathrm{arg}\min}\limits_{\epsilon_{\theta^{\prime}}}\mathbb{E}_{x_{t},t}\left[\lVert\epsilon_{\theta^{\prime}}(x_{t},e_{c},t)-\epsilon_{E}\rVert^{2}\right]$ (10)
$x_{t}\in\mathbb{R}^{W\times H\times d}$ is the denoised image at timestep $t$. $e_{c}$ is the text embedding for concept $c$. $\epsilon_{E}$ is the negatively guided noise predicted by $\theta$, $\epsilon_{E}=\epsilon_{\theta}(x_{t},t)-\eta\left[\epsilon_{\theta}(x_{t},e_{c},t)-\epsilon_{\theta}(x_{t},t)\right]$. By minimizing the distance between $\epsilon_{\theta^{\prime}}(x_{t},e_{c},t)$ and $\epsilon_{E}$, the model learns to lower the probability of the generated image $x$ aligning with the target concept $c$, thereby removing the undesired concept.
Addressing the trade-off between erasing unsafe concepts and preserving locality is a key problem to be solved. Receler [116] trains a lightweight eraser module to erase concepts while ensuring robustness and locality. This is achieved through a concept-localized regularization technique combined with an adversarial prompt learning scheme. EraseDiff [117] frames the unlearning task as a constrained optimization problem, balancing the removal of information related to the data to be forgotten while preserving the model’s performance on the retained data. It accomplishes this by modifying the generative process to diverge from the original denoising procedure associated with the forgotten data. ShieldDiff [130] tackles unsafe content removal in T2I models by fine-tuning a pre-trained diffusion model through reinforcement learning. It utilizes a content-safe reward function specifically designed to produce high-quality, safe images. This function integrates the CLIP model with nudity-specific rewards, enabling effective filtering of inappropriate content while preserving the intended semantics.
Some methods [118, 119, 120, 121] focus on the parameters of the attention module. Forget-Me-Not [118] introduces a plug-and-play method for efficiently forgetting and correcting specific concepts in large-scale text-to-image models. Attention Re-steering is designed to fine-tune only the UNet component to minimize the attention maps corresponding to the target concepts. SalUn [119] introduces the concept of “weight saliency”. By focusing on specific model weights rather than the entire model, SalUn enhances the efficiency and effectiveness of the unlearning process. UCE [120] addresses bias, copyright, and offensive content in text-to-image models by applying edits without additional training. It uses a closed-form solution to modify the model’s attention weights, selectively adjusting the keys and values associated with specific text embeddings for the concepts to be edited while minimizing alterations to the preserved concepts. MACE [121] expands the erasure capacity to handle up to 100 concepts by combining closed-form cross-attention refinement with LoRA fine-tuning, enabling efficient and scalable concept removal.
Several methods [122, 123, 124, 125, 126, 127] use the safe concept and the unsafe concept to guide the model. ESD [122] leverages a brief text description of the unsafe concept to guide the model. It fine-tunes the model by using conditioned and unconditioned scores from the frozen model, steering the output away from the concept being erased. This allows the model to use its internal knowledge to guide the diffusion process, effectively avoiding the undesired concept. SDD [123] applies self-distillation to the diffusion model, guiding the noise estimate conditioned on the target concept to align with the unconditional estimate. It enables the simultaneous removal of multiple concepts in a single process. SafeGen [124] focuses on eliminating explicit unsafe visual representations within the model, regardless of the text input. This approach makes the text-to-image model resistant to adversarial prompts, as it blocks the generation of such unsafe visuals from the inside. SFD [125] encourages the forgetting of undesirable information in diffusion models by aligning the conditional scores of “unsafe” classes or concepts with those of “safe” ones. To avoid relying on real data, the SFD framework integrates a score-based loss into the score distillation objective of a pre-trained diffusion model. UC [126] proposes a domain correction framework using adversarial training to align sensitive and anchor concepts in diffusion models, along with gradient surgery to preserve model performance by mitigating conflicts between unlearning and relearning gradients. Direct Unlearning Optimization (DUO) [127] removes unsafe content from T2I models through preference optimization with paired image data, while output-preserving regularization ensures the model maintains its generative capabilities on safe content.
Different from the previous methods, RECE [128] employs closed-form parameter editing combined with adversarial learning schemes to achieve reliable and efficient concept erasing, ensuring the effective removal of targeted concepts from the model. Another work, Dark Miner [129], follows a recurring three-stage process of mining, verifying, and circumventing. It greedily mines embeddings with the highest generation probabilities for unsafe concepts and effectively reduces the generation of unsafe content by targeting and mitigating these high-risk embeddings.
Guidance-based Defense. Intervening in the forward activations during the inference phase with the external information can guide the model’s output towards safer and more desirable outcomes.
$\bullet$ Image + Text → Text (IT2T) Models. InferAligner[131] proposes to align poorly aligned models for harmlessness by extracting safety steering vectors from pre-aligned models. It targets harmful inputs through intent recognition via safety related vector and utilizes a guidance gate to control activation shifts during inference. By selectively shifting activations in specific transformer layers, InferAligner ensures that only harmful intents are addressed, preserving the model’s performance on other tasks.
$\bullet$ Text → Image (T2I) Models. EIUP[132] introduces an erasure prompt that fuses image features with text embeddings. Through attention mechanisms, it identifies and re-weights non-compliant features, suppressing the generation of unsafe images from implicit unsafe prompts. SLD [93] mitigates inappropriate content generation by combining text conditioning with classifier-free guidance, removing or suppressing unsafe concepts in image outputs. It edits images during inference without fine-tuning, using “unsafe” prompts to guide generation in the opposite direction. Geom-Erasing [133] extracts geometric information of implicit concepts from images, converts it into location tokens, and incorporates these into text prompts. During sampling, these tokens serve as negative prompts, steering the model to generate images free from implicit concepts.
$\bullet$ Text → Video (T2V) Models. SAFREE [134] is a training-free framework that ensures safe T2I and T2V generation by identifying and redirecting prompt token embeddings away from toxic subspaces without altering model weights. It employs a self-validating filtering mechanism to balance toxicity filtering and the preservation of safe concepts, dynamically adjusting denoising steps and using adaptive re-attention within the diffusion latent space. This integration across both textual and visual spaces enables SAFREE to maintain output fidelity, quality, and safety across diverse generative models.
Pruning-based Defense. Unsafe content often originates from concept neurons, which are key to generating specific concepts. Pruning model parameters to deactivate unwanted concept neurons offers an effective strategy for eliminating harmful content.
P-ESD [135] enhances concept-erasing techniques by identifying concept-correlated neurons that are sensitive to adversarial prompts. It selectively removes critical parameters linked to the target concepts to remove undesired concepts. ConceptPrune [136] identifies skilled neurons in feed-forward layers responsible for undesirable concepts in diffusion models.
B.4 Output-level Defense
Output-level Defense leverages post-processing techniques to reinterpret or rephrase the generated content, aiming to neutralize potentially offensive elements while preserving the original intent.
Refiner-based defense. Refiner-based defense typically involves detecting harmful content within generated outputs, followed by employing a generative model to modify the content and eliminate identified toxicity.
$\bullet$ Image + Text → Text (IT2T) Models. MLLM-Protector [137] comprises two core components: a lightweight Harm Detector, which identifies harmful content in model outputs, and a Response Detoxifier, which transforms harmful responses into benign ones. Each component can be independently trained and seamlessly integrated with any MLLM at inference, ensuring robust safety without compromising the model’s original functionality.
$\bullet$ Text → Image (T2I) Models. LMIVS [138] begins with a Visual Commonsense Immorality Recognizer that detects immoral content in generated images. Subsequently, textual and visual immoral attribute localizers identify and highlight specific attributes contributing to the image’s immorality. Once these attributes are identified, the method applies ethical image manipulation techniques to transform the content, ultimately generating morally acceptable outputs.
V Evaluation
Evaluation methods play a crucial role in providing a standardized basis for comparing various jailbreak attack and defense techniques. In this section, we first review existing evaluation datasets related to jailbreak, followed by an overview of metrics used to measure the effectiveness of jailbreak against multimodal generative models.
TABLE V: Comparison of publicly available representative evaluation datasets. For text/images from the data source, 1) Artificial: manually created by humans, 2) Website: scraped from open-world websites, 3) LLM: generated by prompting LLMs, 4) Exist: derived from existing datasets. 5) Adversarial: generated by attack methods, 6) Typography: generated by text typography into image forms, 7) Diffusion: generated by prompting T2I models.
Data Source
Dataset Task Text Image Volume Access
Any-to-Text Models
SafeBench [71] Attack LLM Typography 500 https://github.com/ThuCCSLab/FigStep
AdvBench [54] Attack LLM N/A 500 https://github.com/llm-attacks/llm-attacks
ReadTeam-2K [139] Attack LLM & Exist N/A 2,000 https://huggingface.co/datasets/JailbreakV-28K/JailBreakV-28k
HarmBench [140] Attack & Defense Unpublished N/A 320 https://github.com/centerforaisafety/HarmBench
HADES [86] Defense LLM Typography & Diffusion 750 https://github.com/AoiDragon/HADES
MM-SafetyBench [72] Defense LLM Typography & Diffusion 5,040 https://github.com/isXinLiu/MM-SafetyBench
JailBreakV-28K [139] Defense Adversarial Artificial & Exist & Diffusion 28,000 https://huggingface.co/datasets/JailbreakV-28K/JailBreakV-28k
VLGuard [112] Defense LLM Exist 3,000 https://huggingface.co/datasets/ys-zong/VLGuard
Any-to-Vision Models
NSFW-200 [76] Attack Artificial N/A 200 https://github.com/Yuchen413/text2imagesafety
MMA [19] Attack Exist & Adversarial Adversarial 1,000 https://huggingface.co/YijunYang280
VBCDE-100 [67] Attack Artificial N/A 100 https://github.com/researchcode001/Divide-and-Conquer-Attack
MPUP [141] Attack LLM N/A 1,200 https://huggingface.co/datasets/tongliuphysics/multimodalpragmatic
I2P [93] Attack & Defense Website Website 4,703 https://huggingface.co/datasets/AIML-TUDA/i2p
Unsafe Diffusion [142] Defense Artificial & Website & Exist N/A 1434 https://github.com/YitingQu/unsafe-diffusion
MACE [121] Defense Artificial Diffusion 200 https://github.com/Shilin-LU/MACE
V-A Evaluation Dataset
In this part, we discuss datasets commonly used in jailbreak attacks and defenses for the multimodal generative models. These datasets are categorized into two types based on their construction intent: simulated-malicious and real-malicious datasets. Simulated-malicious datasets involve experiments conducted on pre-existing benchmarks that were not specifically designed for jailbreak scenarios. In contrast, real-malicious datasets are intentionally constructed as representative benchmarks for evaluating jailbreak vulnerabilities in multimodal generative models (see Table V).
V-A1 Simulated-malicious Dataset
Due to the significant malicious potential of unsafe visual content, previous works have sought to mitigate the risk of generating genuinely harmful images in T2I jailbreak attacks. To achieve this, researchers [76, 69, 91] have typically selected specific categories from benign datasets to serve as proxies for unsafe classes, thereby avoiding the generation of real-world malicious imagery during model evaluation and testing. For instance, SneakyPrompt [76] introduces the Dog/Cat-100 dataset, which contains 100 prompts generated by GPT models describing scenarios involving dogs or cats. SneakyPrompt incorporates dog and cat concepts as part of an external safety filter. Similarly, UPAM [69] utilizes the Microsoft COCO dataset [143], a comprehensive resource containing image-text pairs for tasks such as captioning, and designates 10 specific classes (e.g., boat, bird, clock) as ”harmful” categories. Additionally, UnlearnDiffAtk [91] employs the WikiArt dataset [144], which includes fine-art paintings, along with Imagenette [145], to conduct attacks related to stylization and object manipulation.
V-A2 Real-malicious Dataset
Malicious datasets typically contain explicitly harmful prompts, aligned with the usage policies of OpenAI [146] and Meta [147]. These prompts, which often involve content such as pornography, violence, and gore, are widely recognized as critical areas of safety concern due to their potential to cause discomfort and harm. Some commonly used datasets in Table V are as follows:
Image + Text → Text (IT2T) Models. Datasets in this category leverage language models or existing datasets to construct text prompts, with various options available for generating corresponding images. These options include creating blank images, adding noise, typography text onto images, or using Diffusion models to generate malicious images.
$\bullet$ SafeBench [71]. SafeBench identifies 10 safety-critical scenarios and queries GPT-4 with 50 unique malicious questions for each scenario. These questions are then rephrased into imperative sentences and converted into visual prompts through the use of typography.
$\bullet$ AdvBench [54]. AdvBench prompts uncensored LLMs to generate new examples of harmful strings and behaviors by providing 5-shot demonstrations crafted by the authors. The harmful strings represent toxic or harmful topics, while harmful behaviors are framed as instructions that align with the same themes as the harmful strings.
$\bullet$ RedTeam-2K [139]. RedTeam-2K dataset comprises 2,000 meticulously crafted harmful queries aimed at identifying alignment vulnerabilities within MLLMs. It incorporates queries from 16 safety policies and 8 diverse sources, including GPT Rewrite, Handcraft, GPT Generate, and several existing datasets.
$\bullet$ HarmBench [140]. HarmBench encompasses 510 unique harmful behaviors, split into 400 textual behaviors and 110 multimodal behaviors, with categories specifically designed to violate laws or societal norms.
Another line of jailbreak datasets often includes both standard malicious prompts and successful jailbreak prompts, specifically designed to evaluate the vulnerabilities of MLLMs. These datasets provide a rigorous basis for defense evaluation, allowing newly proposed methods to be systematically tested and validated.
$\bullet$ HADES [86]. HADES identified five safety-related scenarios and queried GPT-4 with 50 unique malicious questions for each, pairing these queries with real-world images. Additionally, the datasets include typographic images of harmful keywords and optimized images generated using stable diffusion.
$\bullet$ MM-SafetyBench [72]. MM-SafetyBench queries GPT-4 to generate multiple malicious questions for each scenario and utilizes diffusion models to produce corresponding harmful images based on key phrases extracted from the textual questions. Additionally, the generated images are combined with typography, creating a unified visual representation of the malicious content.
$\bullet$ JailBreakV-28K [139]. Building on the RedTeam-2K dataset, JailBreakV-28K employs a variety of diverse jailbreak attacks to construct jailbreak prompts. These prompts are paired with four types of images: blank images, random noise images, natural images, and synthesized images generated via stable diffusion. The study covers 16 distinct themes and includes a total of 28,000 test cases, consisting of 20,000 text-based jailbreak prompts targeting LLMs and 8,000 image-based jailbreak inputs from recent MLLM jailbreak attacks.
$\bullet$ VLGuard [112]. VLGuard is dedicated to curating an instruction-following dataset that includes various harmful categories. This work gathers harmful images from existing datasets, such as Hateful Memes, and generates a total of 3,000 harmful prompts paired with these images using GPT-4V.
Text → Image (T2I) Models. The datasets for T2I models are primarily focused on evaluating the generation of harmful or inappropriate content through malicious prompts.
$\bullet$ NSFW-200 [76]. NSFW-200 generates 200 target prompts with malicious content using ChatGPT models.
$\bullet$ MMA [19]. MMA consists of 1,000 captions with high harmfulness scores from the LAION-COCO dataset [148]. Additionally, the dataset also incorporates adversarial prompts generated using the corresponding jailbreak methods.
$\bullet$ VBCDE-100 [67]. VBCDE-100 consists of 100 test cases covering five categories of violence, gore, illegal activities, discrimination, and pornographic content, with each category consisting of approximately 20 hand-curated sensitive cues.
$\bullet$ MPUP [141].The MPUP dataset encompasses three high-risk scenarios: hate speech, physical harm, and fraud, with a total of 1,200 unsafe prompts. These prompts are generated by guiding GPT-4 with scenario-specific instructions, followed by manual selection and refinement.
Note that the I2P [93] dataset is widely used as a benchmark for evaluating jailbreak attacks [79, 78, 91, 75, 90] and defenses [132, 127, 128, 124, 110] in T2I models. Attack methods leverage this dataset to construct jailbreak prompts, while defense methods use these prompts to assess the robustness of proposed defense mechanisms against the generated jailbreak inputs.
$\bullet$ I2P [93]. I2P encompasses a broad spectrum of inappropriate content beyond just nudity. This work gathers a total of 4,703 harmful prompts from real-world websites (i.e., Lexica), spanning seven distinct topics, and pairs them with harmful images sourced from those same websites.
In addition, other works have developed their own datasets for research, enabling a more customized approach to studying jailbreak defenses.
$\bullet$ MACE [121]. MACE offers a dataset of 200 celebrities, with their portraits generated through stable diffusion, which the Celebrity Detector identifies with notable accuracy.
$\bullet$ Unsafe Diffusion [142]. Unsafe Diffusion focuses on demystifying the generation of unsafe images and hate memes from T2I models. The study collects harmful prompts from sources such as the Lexica website, and a manually created template-based dataset, along with harmless prompts from the MS COCO dataset, resulting in a collection of 1,434 prompts in total.
Text → Video (T2V) Models. As video generation models continue to evolve, the development of dedicated evaluation datasets for these models has become increasingly crucial.
$\bullet$ T2VSafetyBench [20]. T2VSafetyBench is designed to evaluate the safety of T2V models, focusing on 12 critical safety aspects including pornography, violence, and discrimination. It constructs a dataset of 4,400 malicious prompts from real-world sources, GPT-generated prompts, and various jailbreak attack methods, providing a comprehensive evaluation of model safety across diverse scenarios.
Figure 9: Taxonomy of jailbreak evaluation method against multimodal generative models.
V-B Evaluation Method
In contrast to traditional visual-question answering datasets, generative models produce answers in an open-ended format, complicating subjective evaluation. This open-ended nature presents challenges in balancing evaluation costs with maintaining accuracy. To tackle these challenges, we have outlined diverse evaluation approaches for determining relevant metrics, as illustrated in Fig. 9. These methods are primarily classified into two categories: manual evaluation and automated evaluation.
V-B1 Manual Evaluation
Manual evaluation involves human assessment to determine if the content is toxic, offering a direct and interpretable method of evaluation. Manual evaluation can be categorized into two types: full evaluation and partial evaluation. Full evaluation involves assessing the entire responses, some works [87, 82, 21, 64] have adopted this approach, conducting detailed quantitative experiments to provide a comprehensive analysis of the jailbreak performance. Since full evaluation is labor-intensive and time-consuming, partial dataset evaluation focuses on evaluating a representative subset of the data. For instance, other works [149, 113, 112] use manual evaluation to verify the validity of other automatic evaluations.
V-B2 Automatic Evaluation
Given the high cost and time-consuming nature of human evaluation, recent research has increasingly focused on developing automated evaluation methods. These automated approaches assess the safety of multimodal generative models by employing a range of techniques, including detector-based, GPT-based, and rule-based methods.
TABLE VI: Toxicity detectors are employed as automatic evaluation for multi-modal generative models.
Detector-based. Detector-based approaches utilize pre-trained classifiers to automatically detect and identify harmful content within generated outputs. These classifiers are trained on large, annotated datasets that cover a range of unsafe categories, such as toxicity, violence, or explicit material.
$\bullet$ Text Detector. Toxicity detectors for detecting harmful text encompass several models, including LLama-Guard [150], LLama-Guard2 [151], Detoxify [152], GPTFuzzer [153], and Perspective API [154]. 1) LLama-Guard series models. LLama-Guard, fine-tuned on the LLaMA2 model, functions as a language model capable of performing multi-class classification while also generating binary decision scores for each harmful response. 2) Detoxify. Detoxify is working to mitigate harmful online content by identifying toxic comments. 3) GPTFuzzer. GPTFuzzer identifies jailbreak responses by fine-tuning the RoBERTa model [159] on manually curated datasets with human judgments, enabling it to assess whether a response violates safety protocols. 4) Perspective API. Perspective API employs machine learning models to identify abusive comments by scoring phrases based on their potential impact within a conversation.
$\bullet$ Image Detector. Toxicity detectors for detecting harmful images primarily include NudeNet [155], Q16 [156], SD Safety Checker [157], Imgcensor [158], Multi-headed Safety Classifier (MHSC) [142]. 1) NudeNet. NudeNet detector is employed in the process of categorizing images based on the presence of “nudity”, which has 4 fine-grained nudity labels. 2) Q16. Q16 is fine-tuned using harmful images and corresponding caption data, enabling it to detect a broader range of inappropriate topics, such as violence and misogyny. 3) SD Safety Checker. The SD Safety Checker is integrated into the output stage of Stable Diffusion to detect unsafe elements within generated images. 4) Imgcensor. Imgcensor offers recognition capabilities for images depicting pornography and political figures, utilizing various deep and non-deep learning-based methods. 5) MHSC. MHSC supports fine-grained detection, which builds a multi-headed image safety classifier that detects five unsafe categories simultaneously.
LLM-based. Powerful large language models like GPT-4 have increasingly been employed as safety evaluators in recent research. With carefully crafted prompts, LLMs demonstrate remarkable effectiveness in assessing whether jailbreak behaviors occur, providing nuanced judgments across various scenarios. At evaluation time, LLMs have been employed to evaluate whether the generated content from multimodal large models is semantically consistent with ground-truth annotations [72, 149, 67, 66, 20, 127, 97].
Rule-based. Rule-based approaches generally depend on predefined rules or patterns to assess the alignment between generated content and expected outputs. For instance, keywords can be used to detect the presence of essential concepts, while embedding-based matching ensures that the semantic meaning of the output aligns with ground-truth annotations. These methods offer a structured framework for evaluating both the content accuracy and semantic consistency of generated outputs.
$\bullet$ Keyword-based. Several works [83, 58, 59] on MLLMs have proposed methods to identify responses containing predefined refusal keywords, such as “I am sorry” or “I cannot.” These refusal keywords act as indicators that the model is adhering to safety protocols by refusing to generate harmful or inappropriate content.
$\bullet$ Embedding-based. Other works in T2I tasks [69, 67, 65] utilize embedding-based approaches by comparing the similarity between embeddings generated by CLIP’s encoder. This method ensures that the generated images are semantically aligned with the corresponding semantic concepts by measuring the proximity of their embeddings in the latent space.
V-C Evaluation Metric
V-C1 Metrics for Jailbreak Attack
Evaluating jailbreak attacks on multimodal models requires metrics like Attack Success Rate, which quantifies the proportion of successful attacks that bypass the model’s safety mechanisms.
$\bullet$ Attack Success Rate. The Attack Success Rate (ASR) is a commonly used metric for evaluating the effectiveness of jailbreak methods. It quantifies the proportion of successful attacks out of the total number of attempts. Formally, let $N_{total}$ denotes the total number of jailbreak inputs and $N_{success}$ denotes the number of successful attacks. The ASR can then be defined as:
$ASR=\frac{N_{success}}{N_{total}}$ (11)
$\bullet$ Prompt Perplexity. Perplexity (PPL) [160] serves as a metric for evaluating the readability and linguistic fluency of jailbreak prompts. Many adversarial prompts targeting MLLMs tend to contain garbled or nonsensical characters, which makes them easily detectable and filtered by defense methods that rely on high-perplexity detection. As a result, attack methods that generate low-perplexity prompts are becoming increasingly noteworthy. Formally, let $W=\left(w_{1},w_{2},...,w_{n}\right)$ denote a text sequence where $w_{i}$ represents the $i$-th token. The perplexity of the sequence $W$ is defined as:
$PPL\left(W\right)=\exp\left(-\frac{1}{n}\sum_{i=1}^{n}{\log p_{\theta}\left(w_{i}|w_{<i}\right)}\right)$ (12)
where $p_{\theta}\left(w_{i}|w_{<i}\right)$ represents the probability assigned by MLLMs to the $i$-th token, conditioned on all preceding tokens in the sequence.
V-C2 Metrics for Jailbreak Defense
Evaluating jailbreak defenses in MLLMs often involves monitoring changes in the Attack Success Rate. By comparing ASR before and after the implementation of defense mechanisms against jailbreak prompts, researchers can effectively assess the defenses’ impact. A substantial reduction in ASR suggests that the defense effectively reduces the model’s vulnerability to adversarial inputs, demonstrating its robustness in preventing harmful or unsafe outputs.
For T2I models, Frechet Inception Distance and CLIP Score are utilized to evaluate both the visual quality of generated images and their semantic alignment with adversarial intent following the implementation of defense mechanisms.
$\bullet$ Frechet Inception Distance. Frechet Inception Distance (FID) [161] quantifies the distance between real and generated images. It calculates the distance between the feature representations of real and generated images, as captured by a pre-trained network, thereby assessing how closely the generated images resemble real-world images. The lower the FID score, the better the quality of the generated images, indicating closer alignment with real images. FID is computed as follows:
$FID=\left\|\mu_{r}-\mu_{g}\right\|^{2}+Tr\left(\varSigma_{r}+\varSigma_{g}-2\sqrt{\varSigma_{r}\varSigma_{g}}\right)$ (13)
Where $\mu_{r}$ and $\mu_{g}$ represent the mean vectors of the real and generated image distributions, respectively, and $\varSigma_{r}$ and $\varSigma_{g}$ are the corresponding covariance matrices.
$\bullet$ CLIP Score. CLIP Score [162] measures the alignment of the generated images with their corresponding descriptions. This approach first utilizes the CLIP model $E_{clip}$ to obtain the semantic embeddings of adversarial images. The CLIP Score is then calculated by computing the feature similarity between malicious adversarial images $\mathcal{Y}_{adv}$ and the clean prompts $\mathcal{X}_{clean}$ modified from the toxic content:
$CLIP\,\,Score=Cos\left(E_{clip}\left(\mathcal{Y}_{adv}\right),E_{clip}\left(\mathcal{X}_{clean}\right)\right)$ (14)
VI Future Work
As previously discussed, various studies have investigated jailbreak attacks and defenses for multimodal generative models. However, this area remains in its early stages, with many challenges and potential vulnerabilities yet to be fully explored. This section presents an in-depth analysis of the open challenges and future research directions in multimodal model jailbreaks, aiming to drive further progress toward developing more secure generative systems.
VI-A Multimodal Jailbreak Attack
VI-A1 Underexplored Video/Audio Vulnerabilities
While significant progress has been made in researching jailbreak attacks and defenses for text- and image-based models, video and audio modalities remain underexplored. These modalities are crucial for real-world multi-modal communication, yet their vulnerabilities in generative models, particularly for understanding and generation, are not well understood. The complex temporal dynamics and acoustic features inherent to video and audio pose distinct challenges for evaluating and addressing jailbreak attacks, necessitating more focused research.
VI-A2 Interleaving Multimodal Input
Interleaving multimodal jailbreak attacks pose a significant challenge, as they can simultaneously exploit content across multiple modalities, such as text, images, and audio. While existing research on multimodal jailbreak attacks often involves input from multiple modalities, the attack strategies primarily exploit vulnerabilities in one specific modality, such as text [79, 69, 76, 75] or image [18, 83, 86, 71, 72], rather than addressing the complexities of cross-modal interactions. Therefore, exploring robust techniques for handling interleaving multimodal jailbreak inputs is a crucial future research direction. Conducting interleaving jailbreak attacks across multiple modalities includes the interactions between modalities to simultaneously exploit vulnerabilities in a unified manner. The complexity arises not only from managing multiple data types but also from synchronizing them to achieve a cohesive malicious outcome. This collaborative nature makes detection and defense significantly more difficult, as traditional models are often optimized to handle single-modal inputs.
VI-A3 Diverse Multimodal Output
Attacks targeting Any-to-Any models remain largely unexplored, yet the potential for jailbreak attacks on these models is significant, as they can generate multimodal outputs simultaneously. For instance, a jailbreak attack might involve generating instructions on how to create a bomb, not only producing a textual description, as seen in existing any-to-text models [79, 69, 76, 75, 71, 72], but also generating multiple corresponding images or even sequential content resembling a comic strip. This capability greatly increases the harmful potential of such attacks, as it allows the adversary to convey complex and dangerous information across multiple modalities.
VI-B Multimodal Jailbreak Defense
VI-B1 Developing a Multi-strategy Defense Mechanism
Defense methods based solely on fine-tuning cannot cover a wide range of threats while training-free methods often lack effectiveness and accuracy. Discriminative approaches are also slow to adapt to the evolving strategies of attackers and are therefore incomplete. Future defense strategies can integrate multimodal information to build a robust, layered defense system. This system would combine techniques such as prompt revision, model fine-tuning, and generation guidance to create a more comprehensive and adaptable defense mechanism.
VI-B2 Defense Cost vs. Performance Trade off
Real-time detection of input and generated content introduces significant computational overhead, especially in complex multimodal systems where handling cross-modal information affects inference efficiency. Enhancing defense robustness often comes with high computational costs, potentially degrading model performance. Future research should focus on designing efficient and lightweight defense solutions that maintain strong security without compromising system performance.
VI-B3 Transparency and Explainability of Defense Mechanisms
As multimodal defense mechanisms become more complex, ensuring their transparency and explainability will be a key research focus. Future work should explore ways to make defense systems more transparent, enabling users and developers to understand how these mechanisms function, along with their limitations. Enhancing transparency will improve the credibility and trustworthiness of defense methods.
VI-B4 Research on Personalized Defense Schemes
To address the varying security needs of different users and application scenarios, future research can focus on developing personalized defense schemes. These schemes would allow the strength of defense policies to be dynamically adjusted based on user preferences, specific use cases, and content security levels. For example, stricter defense mechanisms could be activated for sensitive content, while maintaining flexibility and variety for general users.
VI-C Multimodal Jailbreak Evaluation
VI-C1 Limitation in diverse harmful concepts
A key challenge in the evaluation of multimodal models, particularly for T2I systems, is the limited scope of current specialized detectors, which primarily focus on two harmful concepts: violence and pornography. This narrow focus restricts the automated assessment of a broader range of harmful content, such as hate speech, misinformation, and cybercrime. Future work should aim to expand the range of automated detectors to encompass a more comprehensive array of harmful categories. By developing detectors that cover diverse harmful concepts, we can ensure more robust safety evaluations, addressing the complex challenges posed by multimodal models in real-world applications.
VI-C2 Evaluation criteria of multimodal security
Currently, there is no unified evaluation standard in the field of multimodal defense, and the effectiveness of various defense methods cannot be measured systematically. Future work can focus on building a more sophisticated multimodal security evaluation framework to evaluate how defense methods actually perform under different attack scenarios.
VII Conclusion
The increasing prevalence of multimodal generative models has raised significant concerns about their security and reliability, particularly in the context of jailbreak attacks. As a result, extensive research has been devoted to both developing and defending against such attacks. This paper presents a comprehensive overview of jailbreak methods targeting multimodal generative models, systematically categorizing attack and defense techniques across four key levels. We provide a holistic evaluation covering a broad range of modality combinations, including text, image, audio, and video, along with different architectures of the multimodal generative model, such as Any-to-Text, Any-to-Vision, and Any-to-Any models. Additionally, we not only compare existing evaluation datasets, methods, and metrics but also propose critical challenges and potential future directions. While many challenges remain, we hope this paper inspires further discussion and provides strategic guidance for future research, ultimately contributing to the safety and reliability of foundational models.